# Altermagnetic phase transition in a Lieb metal

### v1

@@ Abstract
Introduction—Triggered by the discovery of high-$T_{c}$ superconductivity in copper oxides, the possibility of unconventional pairing transcended our perspective on Fermi surface instabilities. Even though the individual two-electron ground state would always favor zero relative angular momentum according to the Perron-Frobenius theorem [1, 2], the condensation energy gain to the quantum fluid formed by phase-coherent Cooper pairs drives the preference of a $d$-wave particle-particle condensate, i.e., a pairing state with finite relative angular momentum. While the field has been aware of its principal possibility for decades, it has proven challenging to microscopically identify a similarly unconventional magnet, i.e., a spinful particle-hole condensate with finite relative angular momentum [3]. In combination with translation symmetry breaking, a rare instance is the kagome Hubbard model exhibiting a $p$-wave spin bond order [4]. In combination with point group symmetry breaking, a spin-type $d$-wave nematic order has been claimed in the context of higher order van-Hove singularities [5], but has not yet unfolded in a generic microscopic model. The omnipresent difficulty in realizing such states is engraved in the oppositely charged constituents of the particle-hole pairs forming the quantum fluid, as they are naturally expected to gain condensation energy from zero relative angular momentum between particle and hole [6, 7].
@@
A new chapter on broadening the phenomenological scope of exotic magnetic order has started with the discovery of a seemingly unprecedented type of magnetic order in 2019, now termed altermagnetism (AM) [8, 9, 10, 11, 12, 13]. The key driving interest rests upon the potential use of AMs for spintronics, as their energetically spin-split magnetic excitations, along with zero net magnetization, advantageously combine antiferromagnetic and ferromagnetic features for technological utilization [14, 15, 16, 17]. A collinear AM is designed such that sites with opposite-spin orientation transform into each other by spatial rotation [18, 19]. While this conceptually resembles the intertwining of time reversal and point group symmetry related to spin Pomeranchuk instabilities, a core insight embodied by the rising AM research domain is that such intertwining can just as well be more trivially achieved in a scale-separated cascade of transitions: As of now, the aspired abundance of AMs builds on the formation of anisotropies at the crystal field energy scale, followed by an independent magnetic transition determined by local moments and additional low-energy electronic scales [20, 21]. A complete magnetic analogue of unconventional superconducting pairing, however, would only be resembled by the microscopic realization of a metallic parent state subject to a single phase transition into AM order [22, 23].
@@
In this Letter, we formulate a microscopic realization of an altermagnetic phase transition for interacting electrons on the Lieb lattice. Previous attempts to overcome the scale-separated nature of AM formation revolved around interacting electrons subject to combined staggered orbital and antiferromagnetic ordering. These proposals, however, are contrived in some aspects, as the natural propensity of a staggered orbitally ordered system to form a ferromagnet due to Hund’s coupling can only be avoided by ignoring the Hund’s coupling altogether [24, 25]. Instead, we seek to intertwine rotation and time-reversal symmetry through the sublattice interference (SI) profile found in a single-orbital Lieb metal parent state [26]. Originally introduced for the kagome lattice [27, 4, 28], SI ascribes pivotal relevance to the sublattice distribution of Fermi level eigenstates with regard to identifying the relevant scattering channels for the unfolding electronic order. On the Lieb lattice, such interference leads to a dissociation of sublattice contributions to the magnetic fluctuation profile of the Lieb metal parent state. Through functional renormalization group (FRG) calculations [29, 30, 31, 32], we show how this yields a phase transition into an AM phase. We classify the emerging order to descend from a $d$-wave spin Pomeranchuk instability, and to reveal a spin-split quasiparticle bandstructure featuring symmetry-protected nodal lines.
@@
Figure 1: (a) Real space lattice structure with of the Lieb lattice. The purple square marks the unit cell with the $A$ site on the trivial Wyckoff position ($1a$), and the $B$ and $C$ sites on the $2c$ Wyckoff positions. Nearest- and next-nearest neighbor hoppings $t$ and $t^{\prime}$ are indicated. The gray square corresponds to the unit cell of a lattice without the $A$ site, i.e., a simple square lattice. (b) Resulting band structure along the high-symmetry path indicated in panel (c) for $t^{\prime}=t/2$ and $\mu_{A}=0$. We overlay the sublattice content by color (purple: $A$, gray: $B$/$C$). The right panel displays the sublattice resolved density of states, with the $B$/$C$ polarized van-Hove singularity at the band touching point. (c) Brillouin zones (BZ) of the Lieb lattice structure (purple) and the intercalated square lattice (gray) formed by dashed lines connecting the $2c$ Wyckoff positions. We additionally plot the Fermi surface and its sublattice polarization (yellow: $B$, green: $C$), which is pure along the BZ boundaries and mixed at the van-Hove points $M$.
@@
Model—We start from tightly bound electrons on a Lieb lattice, which is a square-depleted 2D lattice with three sites per unit cell and features equivalent nearest neighbor bonds connecting inequivalent sites. While it is rarely the lattice structure that minimizes a generic crystal field potential, the Lieb lattice forms a central building block of (layered) perovskites [33], and as such is relevant for high-$T_{c}$ copper oxides when the oxygen sites are not integrated out, but kept for their low-energy description [34]. Figure 1a illustrates the sublattices on Wyckoff positions $1a$ ($A$) and $2c$ ($B$, $C$). While the $A$ site transforms trivially under all elements of $C_{4v}$, the sites $B$ and $C$ map onto each other under $C_{4}$ rotations. This transformation behavior is inherited by the tight-binding Hamiltonian. It reads
@@
where the operator $c^{({{\dagger}})}_{i\sigma}$ annihilates (creates) an electron with spin $\sigma$ on site $i$, $n_{i\sigma}=c^{{\dagger}}_{i\sigma}c^{\vphantom{{\dagger}}}_{i\sigma}$ is the density operator, $\mu_{i}$ denotes the chemical potential at site $i$, and $U$ is the Hubbard onsite repulsion. We employ an extended Lieb lattice model by setting $t_{ij}=1$ for nearest neighbors (NN, $A$-$B$ and $A$-$C$ bonds, solid lines in Fig. 1a) and $t_{ij}=t^{\prime}$ for next-nearest neighbors (NNN, $B$-$C$ bonds, dashed lines). To allow for the effect of different atomic species at Wyckoff positions $1a$ ($A$) and $2c$ ($B$, $C$), we include an intrinsic detuning $\mu_{A}\neq\mu_{B,C}$.
@@
While the electronic spectrum of the metallic parent state displays the full space group symmetry, the multi-site unit cell exhibits a non-trivial transformation behaviour of the eigenstates (Bloch functions). For the Lieb lattice with NN hopping only ($t^{\prime}=0$), a flat band resides at the Fermi level featuring a distinct sublattice polarization on the zone boundary: Excluding the zone corner at $M$, only the $B$ and $C$ sites contribute to the electronic eigenstates (see SM [35] for details) [26]. While this feature persists in the presence of NNN hopping $t^{\prime}$, the flat band attains a sizable dispersion away from the zone boundary, producing a van-Hove singularity (VHS, see Fig. 1b). At the van-Hove $M$ point, the system displays a topologically protected quadratic (triple) band touching for $\mu_{A}\neq 0$ ($\mu_{A}=0$[33]. Even at pristine filling (dashed line in Fig. 1b), the proximate VHS implies a large density of states (DOS) participating in a tentative Fermi surface instability, rendering the system already unstable against weak repulsive interactions. The single-particle spectrum and sublattice polarization can be understood by decomposing the extended Lieb lattice into two intercalated square lattices. The $B$ and $C$ sublattice form a regular square lattice with lattice constant $1/\sqrt{2}$ and NN hopping $t^{\prime}$, which is coupled to the larger square lattice of the site $A$ via $t$ (cf. Fig. 1a). Comparing the Brillouin zone (BZ) of the associated unit cells in Fig. 1c, the Fermi surface (FS) at the zone boundary reveals itself as the characteristic FS of the smaller square lattice (gray) backfolded to the Lieb lattice BZ (purple). While the weight of the $B$ and $C$ sublattices must be equal and constant in the the $t=0$ limit, the Lieb lattice structure allows for SI along the FS, manifesting in $B$/$C$ polarization along different directions. Regardless of the value of $t^{\prime}$, we find that the eigenstate along $X$-$M$ ($Y$-$M$) displays full $C$ ($B$) polarization. In turn, the high-energy VHSs display a strong $AB/AC$ mixing (see Fig. 1b). This behavior is reminiscent of the pure/mixed VHSs found in the kagome lattice [36, 28, 37, 38].
@@
Figure 2: $d$-wave altermagnetic state on the Lieb lattice. (a) Real space magnetization pattern in the $B_{1}$ irreducible representation of $C_{4v}$. The magnetic order parameter $\Delta$ is nonzero only on the $B$ and $C$ sites. (b) Quasiparticle bandstructure in the altermagnetic phase for $t^{\prime}=t/2$, $\mu_{A}=0$, and $\Delta_{\mathrm{M}}=0.2\,t$ with spin polarization indicated by purple/green ($\downarrow$/$\uparrow$) lines. The inset shows the spin polarized Fermi surface, where the nontrivial transformation behavior under $C_{4}$ rotations and diagonal mirrors $\mathcal{M}_{xy}$, $\mathcal{M}_{x\bar{y}}$ becomes apparent.
@@
In the Lieb lattice model (cf. Fig. 1a), we choose the Fermi level at the VHS close to half filling, expecting ordering tendencies to be strongest at points of high DOS. Given the FS geometry shown in Fig. 1c, the FRG flows towards a magnetic instability irrespective of minor changes to the model [35]. The renormalized interaction is dominated by particle-hole nesting contributions with transfer momentum $\bm{q}=\bm{g}_{1,2}\equiv\Gamma$, with $\bm{g}_{1,2}$ reciprocal lattice vectors. Such “ferromagnetic” (disregarding sublattice structure) fluctuations have previously been ascribed to the Lieb lattice’s non-trivial quantum-geometry [45], which is captured by TUFRG. Figure 2a presents a real-space picture of the magnetic order parameter $\Delta$, which has been found to be stable at the mean-field level [33]. In momentum space, its $d$-wave transformation behavior ($B_{1}$ irreducible representation of $C_{4v}$) becomes apparent, i.e.,
@@
with $\Delta_{M}$ representing the magnitude and $\hat{\sigma}_{z}$ the Pauli-$z$ matrix in spin space. While the $B$/$C$ sites feature a sizable spin polarization, the $A$ site does not participate in the magnetic ordering process. The continuous deformation of the Fermi surface at the phase transition (cf. Fig. 2b) in the absence of translation symmetry breaking classifies the state as an $l=2$ spin Pomeranchuk instability (sPI) [22, 23]. Compared to the original proposal of Pomeranchuk, the Fermi momenta themselves do not exhibit a reduced symmetry; rather, only the spin polarization along the FS acquires a dependence on the Fermi momentum. This implies TRS breaking without Kramers degeneracy, and provides the non-relativistic $P$-$2$ spin momentum locking characteristic for $d$-wave AMs [21].
In the Lieb lattice structure, the absence of a finite magnetisation on the $A$ site is tightly related to the emergence of an AM state [39]. If a site on the trivial Wyckoff position $1a$ contributes to the magnetic order, the intra-unit cell structure of the magnetic ordering vector is bound to the trivial $A_{1}$ representation, and a spin-compensated magnetic state can thus only arise from broken translation symmetry, which would result in an AFM-type state. Intuitively, a finite magnetisation on the $A$ site might be expected due to the higher coordination of the $1a$ Wyckoff position, and hence the resulting increased energy gain in the ordered state. Instead, the itinerant nature of the present instability counteracts a gap on the $A$ site by SI. In particular around van-Hove filling, the scattering channels involve only the $B$ and $C$ sites. Notably, magnetization patterns involving the $A$ site are suppressed for $t>t^{\prime}$ (i.e., the materials-oriented setting, see SM [35]), promoting AM order on the $2c$ positions  [^1] In the limit $t^{\prime}=0$, Lieb’s theorem holds, and we expect a finite magnetization. Setting $t^{\prime}\neq 0$, however, does not allow for the same conclusion, enabling AM order to emerge.. In more general terms, the emergence of the AM state is tied to the sublattice structure at the VHS, which persists even in the case of longer-ranged hoppings and interactions [35]. Our results highlight that in the itinerant picture, magnetic ordering on the $A$ site is not suppressed a priori by crystal field effects. Instead, the precise nature electronic band structure states at the Fermi level enables SI to percolate into the renormalized interaction, and eventually to favor an AM instability.
@@
Figure 2b demonstrates how the quasiparticle band structure inherits the non-trivial transformation behavior of the $2c$ sites under $C_{4}$ rotations and diagonal mirrors $\mathcal{M}_{xy}$, $\mathcal{M}_{x\bar{y}}$: In addition to the non-relativistic spin-splitting along $\Gamma$-$X$ ($\Gamma$-$Y$), the magnetic gap $\Delta(\bm{k})$ exhibits symmetry-protected nodal lines along $\Gamma$-$M$. To gain insights on the AM phase transition, we use our FRG result to perform a constrained self-consistent mean-field analysis [47, 48, 35]. We find a second-order transition with $\Delta$ smoothly increasing from $T=T_{c}$ to its maximum value at $T=0$ (cf. Fig. 3a). At the VHS, where the divergent DOS leads to the largest induced spin gap, we demonstrate that significant values of $\Delta$ can be reached: Figure 3b displays the mean-field result for $\Delta_{M}$ as a function of $U$ at zero temperature. Even in the weak to intermediate coupling regime ($U/W<1$, white region), $\Delta_{M}$ is on the order of $t$. In turn, the size of the spin splitting $\delta$ is readily given by the magnitude of the altermagnetic order parameter $\Delta_{M}$ for $\Delta_{M}\ll t$ [35] to read
$\delta(\bm{k}=X)=\begin{cases}2\Delta_{M}&\quad\text{for the central band,}\\ \Delta_{M}\,|1\pm\mu_{A}/t|&\quad\text{for the outer bands.}\end{cases}$ (3)
@@
Discussion—We show how an itinerant electronic Fermi surface instability mechanism can lead to an altermagnetic phase transition. Instead of the conventional hierarchy in altermagnetic materials, where a high-energy crystal field effect precedes magnetic ordering at lower energy scales, our approach makes use of the sublattice texture imprinted into electronic correlations through the Fermi level. As a consequence, only some of the degrees of freedom are selected in the formation of a magnetic state, eventually leading to a spin Pomeranchuk instability. In the ordered phase, the quasiparticle band structure displays features known from altermagnets, i.e., non-relativistic spin splitting with net zero magnetization in the absence of translation symmetry breaking. We illustrate the itinerant mechanism through the example of a Hubbard model on the Lieb lattice. By means of the functional renormalization group, we obtain its weak coupling instabilities and find a prominent tendency towards an $l=2$ ($d$-wave) spin Pomeranchuk order. Notably, the crystal structure does not fall into the classification of altermagnets [49, 50, 51, 39]. Instead, sublattice interference leads to selective site participation in the magnetic order. Our obtained magnetization pattern then is equivalent to the $P$-$2$ altermagnetic state featured in toy models of RuO2 [20, 39].
The absence of altermagnetic signatures in RuO2 [52, 53] as well as MnF2 [54] candidate materials is remarkable in face of an expectation of large AM energy scale from mean field theory. This might point to a substantial screening effect so far overlooked in the hierarchical, scale-separated symmetry breaking mechanism of local moment AM setups, deserving further investigation. In addition, Ref. [55] demonstrates that altermagnetic metals are difficult to find assuming hierarchical symmetry breaking. In this context, the lack of scale separation embodied by our AM mechanism establishes a notable difference, which might bring about beneficial conditions for AM order and its descendant spin splitting. We expect our itinerant AM formation principle to unlock new altermagnetic states in crystallographic geometries outside the conventional classification of altermagnets presented in Refs. [49, 50, 51, 39], given the abundance of sublattice interference in various two- and three-dimensional lattices [26]. Since our AM mechanism offers a direct transition from a metallic to an altermagnetic state, it can be probed by, e.g., spin-dependent scanning tunneling microscopy or angle-resolved photoemission spectroscopy. Some candidate systems include covalent organic systems [56], metal-organic frameworks [57], and optical lattices [35]. Furthermore, the approach presented here is not restricted to collinear AM order, but can likewise be applied in the context of non-collinear altermagnetism [58].
@@
S2 Effective model for the $B$/$C$ sites
@@
$\mathcal{G}_{BC}=\frac{1}{{\big{(}{\mathcal{G}}_{BC}^{0}\big{)}}^{-1}-\frac{T{{\mathcal{G}}_{A}^{0}}T^{\dagger}}{{\big{(}{\mathcal{G}}_{BC}^{0}\big{)}}^{-1}-\frac{T{{\mathcal{G}}_{A}^{0}}T^{\dagger}}{{\big{(}{\mathcal{G}}_{BC}^{0}\big{)}}^{-1}-\dots}}}\approx\frac{1}{i\omega-H_{BC}-T{{\mathcal{G}}_{A}^{0}}T^{\dagger}}\equiv\frac{1}{i\omega-H_{BC}-\Sigma_{BC}}\,,$ (S5)
where the approximation corresponds to a truncation of the continued fraction at leading order. The (hybridization) self-energy $\Sigma_{BC}$ is therefore given as

### v2

@@ Abstract
Introduction—Triggered by the discovery of high-$T_{c}$ superconductivity in copper oxides, the possibility of unconventional pairing transcended our perspective on Fermi surface instabilities. Even though the individual two-electron ground state would always favor zero relative angular momentum according to the Perron-Frobenius theorem [55, 19, 66, 30], the condensation energy gain to the quantum fluid formed by phase-coherent Cooper pairs drives the preference of a $d$-wave particle-particle condensate, i.e., a pairing state with finite relative angular momentum. While the field has been aware of its principal possibility for decades, it has proven challenging to microscopically identify a similarly unconventional magnet, i.e., a spinful particle-hole condensate with finite relative angular momentum [80]. In combination with translation symmetry breaking, a rare instance is the kagome Hubbard model exhibiting a $p$-wave spin bond order [35]. In combination with point group symmetry breaking, a spin-type $d$-wave nematic order has been claimed in the context of higher order van-Hove singularities [10], but has not yet unfolded in a generic microscopic model. The omnipresent difficulty in realizing such states is engraved in the oppositely charged constituents of the particle-hole pairs forming the quantum fluid, as they are naturally expected to gain condensation energy from zero relative angular momentum between particle and hole [37, 82].
@@
A new chapter on broadening the phenomenological scope of exotic magnetic order has started with the discovery of a seemingly unprecedented type of magnetic order in 2019, now termed altermagnetism (AM)  [69, 52, 2, 1, 83, 65, 22, 18, 53, 84, 25, 46, 48, 75]. The key driving interest rests upon the potential use of AMs for spintronics, as their energetically spin-split magnetic excitations, along with zero net magnetization, advantageously combine antiferromagnetic and ferromagnetic features for technological utilization [21, 67, 70, 24]. A collinear AM is designed such that sites with opposite-spin orientation transform into each other by spatial rotation [85, 7]. While this conceptually resembles the intertwining of time reversal and point group symmetry related to spin Pomeranchuk instabilities, a core insight embodied by the rising AM research domain is that such intertwining can just as well be more trivially achieved in a scale-separated cascade of transitions: A few proposals start from the formation of magnetic moments at high energy combined with a subsequent rotation symmetry breaking at lower energies [5, 47, 4, 86]. As of now, the aspired abundance of AMs, however, mainly builds on the formation of anisotropies at the crystal field energy scale, followed by an independent magnetic transition determined by local moments and additional low-energy electronic scales [72, 71]. A complete magnetic analogue of unconventional superconducting pairing, however, would only be resembled by the microscopic realization of a metallic parent state subject to a single phase transition into AM order [58, 9].
@@
In this Letter, we formulate a microscopic realization of an altermagnetic phase transition for interacting electrons on the Lieb lattice. Previous attempts to overcome the scale-separated nature of AM formation revolved around interacting electrons subject to combined staggered orbital and antiferromagnetic ordering. These proposals, however, are contrived in some aspects. In particular, the natural propensity of a staggered orbitally ordered system to form a ferromagnet due to Hund’s coupling can only be avoided by ignoring the Hund’s coupling altogether [40, 20]. Instead, we seek to intertwine rotation and time-reversal symmetry through the sublattice interference (SI) profile found in a single-orbital Lieb metal parent state [43]. Originally introduced for the kagome lattice [36, 35, 31], SI ascribes pivotal relevance to the sublattice distribution of Fermi level eigenstates with regard to identifying the relevant scattering channels for the unfolding electronic order. On the Lieb lattice, such interference leads to a dissociation of sublattice contributions to the magnetic fluctuation profile of the Lieb metal parent state. Through functional renormalization group (FRG) calculations [64, 50, 56, 15], we show how this yields a phase transition into an AM phase. We classify the emerging order to descend from a $d$-wave spin Pomeranchuk instability, and to reveal a spin-split quasiparticle bandstructure featuring symmetry-protected nodal lines.
@@
Figure 1: (a) Real space lattice structure with of the Lieb lattice. The purple square marks the unit cell with the $A$ site on the trivial Wyckoff position ($1a$), and the $B$ and $C$ sites on the $2c$ Wyckoff positions. Nearest- and next-nearest neighbor hoppings $t$ and $t^{\prime}$ are indicated. The gray square corresponds to the unit cell of a lattice without the AA site, i.e., a simple square lattice. (b) Resulting band structure along the high-symmetry path indicated in panel (c) for $t^{\prime}=t/2$ and $\mu_{A}=0$. We overlay the sublattice content by color (purple: AA, gray: BB/$C$). The right panel displays the sublattice resolved density of states, with the BB/$C$ polarized van-Hove singularity at the band touching point. (c) Brillouin zones (BZ) of the Lieb lattice structure (purple) and the intercalated square lattice (gray) formed by dashed lines connecting the $2c$ Wyckoff positions. We additionally plot the Fermi surface and its sublattice polarization (yellow: $B$, green: $C$), which is pure along the BZ boundaries and mixed at the van-Hove points $M$.
@@
Model—We start from tightly bound electrons on a Lieb lattice, which is a square-depleted 2D lattice with three sites per unit cell and features equivalent nearest neighbor bonds connecting inequivalent sites. While it is rarely the lattice structure that minimizes a generic crystal field potential, the Lieb lattice forms a central building block of (layered) perovskites [76], and as such is relevant for high-$T_{c}$ copper oxides when the oxygen sites are not integrated out, but kept for their low-energy description [17]. Figure 1a illustrates the sublattices on Wyckoff positions $1a$ (AA) and $2c$ ($B$, $C$). While the $A$ site transforms trivially under all elements of $C_{4v}$, the sites $B$ and $C$ map onto each other under $C_{4}$ rotations. This transformation behavior is inherited by the tight-binding Hamiltonian. It reads
@@
where the operator $c^{({{\dagger}})}_{i\sigma}$ annihilates (creates) an electron with spin $\sigma$ on site $i$, $n_{i\sigma}=c^{{\dagger}}_{i\sigma}c^{\vphantom{{\dagger}}}_{i\sigma}$ is the density operator, $\mu_{i}$ denotes the chemical potential at site $i$, and $U$ is the Hubbard onsite repulsion. We employ an extended Lieb lattice model by setting $t_{ij}=1$ for nearest neighbors (NN, AA-$B$ and AA-$C$ bonds, solid lines in Fig. 1a) and $t_{ij}=t^{\prime}$ for next-nearest neighbors (NNN, $B$-$C$ bonds, dashed lines). To allow for the effect of different atomic species at Wyckoff positions $1a$ (AA) and $2c$ ($B$, $C$), we include an intrinsic detuning $\mu_{A}\neq\mu_{B,C}$.
@@
While the electronic spectrum of the metallic parent state displays the full space group symmetry, the multi-site unit cell exhibits a non-trivial transformation behaviour of the eigenstates (Bloch functions). For the Lieb lattice with NN hopping only ($t^{\prime}=0$), a flat band resides at the Fermi level featuring a distinct sublattice polarization on the zone boundary: Excluding the zone corner at $M$, only the $B$ and $C$ sites contribute to the electronic eigenstates (see SM [74] for details) [43]. While this feature persists in the presence of NNN hopping $t^{\prime}$, the flat band attains a sizable dispersion away from the zone boundary, producing a van-Hove singularity (VHS, see Fig. 1b). At the van-Hove $M$ point, the system displays a topologically protected quadratic (triple) band touching for $\mu_{A}\neq 0$ ($\mu_{A}=0$[76]. Even at pristine filling (dashed line in Fig. 1b), the proximate VHS implies a large density of states (DOS) participating in a tentative Fermi surface instability, rendering the system already unstable against weak repulsive interactions. The single-particle spectrum and sublattice polarization can be understood by decomposing the extended Lieb lattice into two intercalated square lattices. The $B$ and $C$ sublattice form a regular square lattice with lattice constant $1/\sqrt{2}$ and NN hopping $t^{\prime}$, which is coupled to the larger square lattice of the site $A$ via $t$ (cf. Fig. 1a). Comparing the Brillouin zone (BZ) of the associated unit cells in Fig. 1c, the Fermi surface (FS) at the zone boundary reveals itself as the characteristic FS of the smaller square lattice (gray) backfolded to the Lieb lattice BZ (purple). While the weight of the $B$ and $C$ sublattices must be equal and constant in the the $t=0$ limit, the Lieb lattice structure allows for SI along the FS, manifesting in BB/$C$ polarization along different directions. Regardless of the value of $t^{\prime}$, we find that the eigenstate along $X$-$M$ ($Y$-$M$) displays full $C$ ($B$) polarization. In turn, the high-energy VHSs display a strong $AB/AC$ mixing (see Fig. 1b). This behavior is reminiscent of the pure/mixed VHSs found in the kagome lattice [81, 31, 28, 33].
@@
Figure 2: $d$-wave altermagnetic state on the Lieb lattice. (a) Real space magnetization pattern in the $B_{1}$ irreducible representation of $C_{4v}$. The magnetic order parameter $\Delta$ is nonzero only on the $B$ and $C$ sites. (b) Quasiparticle bandstructure in the altermagnetic phase for $t^{\prime}=t/2$, $\mu_{A}=0$, and $\Delta_{\mathrm{M}}=0.2\,t$ with spin polarization indicated by purple/green ($\downarrow$$\uparrow$) lines. The inset shows the spin polarized Fermi surface, where the nontrivial transformation behavior under $C_{4}$ rotations and diagonal mirrors $\mathcal{M}_{xy}$, $\mathcal{M}_{x\bar{y}}$ becomes apparent.
@@
In the Lieb lattice model (cf. Fig. 1a), we choose the Fermi level at the VHS close to half filling, expecting ordering tendencies to be strongest at points of high DOS. Given the FS geometry shown in Fig. 1c, the FRG flows towards a magnetic instability irrespective of minor changes to the model [74]. The renormalized interaction is dominated by commensurate particle-hole nesting contributions with transfer momentum $\bm{q}=\bm{g}_{1,2}\equiv\Gamma$, with $\bm{g}_{1,2}$ reciprocal lattice vectors see Section A of the SM [74] for details on interpreting TUFRG results). Such “ferromagnetic” (disregarding sublattice structure) fluctuations have previously been ascribed to the Lieb lattice’s non-trivial quantum-geometry [38], which is captured by TUFRG. Figure 2a presents a real-space picture of the magnetic order parameter $\Delta$, which has been found to be stable at the mean-field level [76]. In momentum space, its $d$-wave transformation behavior ($B_{1}$ irreducible representation of $C_{4v}$) becomes apparent, i.e.,
@@
with $\Delta_{M}$ representing the magnitude and $\hat{\sigma}_{z}$ the Pauli-$z$ matrix in spin space. While the BB/$C$ sites feature a sizable spin polarization, the $A$ site does not participate in the magnetic ordering process. The continuous deformation of the Fermi surface at the phase transition (cf. Fig. 2b) in the absence of translation symmetry breaking classifies the state as an $l=2$ spin Pomeranchuk instability (sPI) [58, 9]. Compared to Pomeranchuk’s proposal, the Fermi momenta themselves do not exhibit a reduced symmetry; rather, only the spin polarization along the FS acquires a dependence on the Fermi momentum. This implies TRS breaking without Kramers degeneracy, and provides the non-relativistic PP-22 spin momentum locking characteristic for $d$-wave AMs [71].
In the Lieb lattice structure, the absence of a finite magnetisation on the $A$ site is tightly related to the emergence of an AM state [63]. If a site on the trivial Wyckoff position $1a$ contributes to the magnetic order, the intra-unit cell structure of the magnetic ordering vector is bound to the trivial $A_{1}$ representation, and a spin-compensated magnetic state can thus only arise from broken translation symmetry, which would result in an AFM-type state. Intuitively, a finite magnetisation on the $A$ site might be expected due to the higher coordination of the $1a$ Wyckoff position, and hence the resulting increased energy gain in the ordered state. Instead, the itinerant nature of the present instability counteracts a gap on the $A$ site by SI. In particular around van-Hove filling, the scattering channels involve only the $B$ and $C$ sites. Notably, magnetization patterns involving the $A$ site are suppressed for $t>t^{\prime}$ (i.e., the materials-oriented setting, see SM [74]), promoting AM order on the $2c$ positions  [^1] In the limit $t^{\prime}=0$, Lieb’s theorem holds, and we expect a finite magnetization. Setting $t^{\prime}\neq 0$, however, does not allow for the same conclusion, enabling AM order to emerge.. In more general terms, the emergence of the AM state is tied to the sublattice structure at the VHS, which persists even in the case of longer-ranged hoppings and interactions [74]. Our results highlight that in the itinerant picture, magnetic ordering on the $A$ site is not suppressed a priori by crystal field effects. Instead, the precise nature of the states at the Fermi level enables SI to percolate into the renormalized interaction, and eventually to favor an AM instability.
@@
Figure 2b demonstrates how the quasiparticle band structure inherits the non-trivial transformation behavior of the $2c$ sites under $C_{4}$ rotations and diagonal mirrors $\mathcal{M}_{xy}$, $\mathcal{M}_{x\bar{y}}$: In addition to the non-relativistic spin-splitting along $\Gamma$-$X$ ($\Gamma$-$Y$), the magnetic gap $\Delta(\bm{k})$ exhibits symmetry-protected nodal lines along $\Gamma$-$M$. To gain insights on the AM phase transition, we use our FRG result to perform a constrained self-consistent mean-field analysis (see Section C of the SM [74] and Refs. [78, 54]). We find a second-order transition with $\Delta$ smoothly increasing from $T=T_{c}$ to its maximum value at $T=0$ (cf. Fig. 3a). At the VHS, where the divergent DOS leads to the largest induced spin gap, we demonstrate that significant values of $\Delta$ can be reached: Figure 3b displays the mean-field result for $\Delta_{M}$ as a function of $U$ at zero temperature. Even in the weak to intermediate coupling regime ($U/W<1$, white region), $\Delta_{M}$ is on the order of $t$. In turn, the size of the spin splitting $\delta$ is readily given by the magnitude of the altermagnetic order parameter $\Delta_{M}$ for $\Delta_{M}\ll t$ [74] to read
$\delta|_{\bm{k}=X}=\begin{cases}2\Delta_{M}&\quad\text{for the central band,}\\ \Delta_{M}\,|1\pm\mu_{A}/t|&\quad\text{for the outer bands.}\end{cases}$ (3)
@@
Discussion—We show how an itinerant electronic Fermi surface instability mechanism can lead to an altermagnetic phase transition. Instead of the conventional hierarchy in altermagnetic materials, where a high-energy crystal field effect precedes magnetic ordering at lower energy scales, our approach makes use of the sublattice texture imprinted into electronic correlations through the Fermi level. As a consequence, only some of the degrees of freedom are selected in the formation of a magnetic state, leading up to a spin Pomeranchuk instability. In the ordered phase, the quasiparticle band structure displays features known from altermagnets, i.e., non-relativistic spin splitting with net zero magnetization in the absence of translation symmetry breaking. We substantiate the itinerant mechanism at the example of a Hubbard model on the Lieb lattice. By means of the FRG, we obtain its weak coupling instabilities and find a prominent tendency towards an $l=2$ ($d$-wave) spin Pomeranchuk order. Notably, the crystal structure does not fall into the classification of altermagnets [23, 79, 73, 63]. Instead, sublattice interference leads to selective site participation in the magnetic order. Our obtained magnetization pattern then is equivalent to the PP-22 altermagnetic state featured in toy models of RuO2 [72, 63].
The absence of altermagnetic signatures in RuO2 [44, 57] as well as MnF2 [51] candidate materials is remarkable in face of an expectation of large AM energy scale from mean field theory. This might point to a substantial screening effect so far overlooked in the hierarchical, scale-separated symmetry breaking mechanism of local moment AM setups, deserving further investigation. In addition, Ref. [77] suggests that altermagnetic metals are rather challenging to find assuming hierarchical symmetry breaking. In this context, the lack of scale separation embodied by our AM mechanism establishes a notable difference, which might bring about beneficial conditions for AM order and its descendant spin splitting. We expect our itinerant AM formation principle to unlock new altermagnetic states in crystallographic geometries outside the conventional classification of altermagnets presented in Refs. [23, 79, 73, 63], given the abundance of sublattice interference in various two- and three-dimensional lattices [43]. Since our AM mechanism offers a direct transition from a metallic to an altermagnetic state, it can be probed by, e.g., spin-dependent scanning tunneling microscopy or angle-resolved photoemission spectroscopy. Some candidate systems include covalent organic systems [11], metal-organic frameworks [6], and optical lattices (see SM [74] for the Lieb lattice, and Ref. [12] for a square lattice proposal). Furthermore, the approach presented here is not restricted to collinear AM order, but can likewise be applied in the context of non-collinear altermagnetism [8].
@@
S2 Effective model for the BB/$C$ sites
@@
$\mathcal{G}_{BC}=\frac{1}{{\big({\mathcal{G}}_{BC}^{0}\big)}^{-1}-T{{\mathcal{G}}_{A}^{0}}T^{\dagger}}={}\frac{1}{i\omega-H_{BC}-T{{\mathcal{G}}_{A}^{0}}T^{\dagger}}\equiv\frac{1}{i\omega-H_{BC}-\Sigma_{BC}}\,.$ (S5)
The (hybridization) self-energy $\Sigma_{BC}$ is therefore given as